OpenStudy (beccaboo333):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and second equations, then the second and third equations. 3x - 5y + 2z = 7 4x + 9y - z = 2 x - 3y + 3z = 7

OpenStudy (ranga):

To eliminate z from the first two equations, multiply the second equation by 2 and add it to the first. To eliminate z from the last two equations, multiply the second equation by 3 and add it to the third. What you will be left with are two equations and two unknowns.

OpenStudy (beccaboo333):

O.o

OpenStudy (anonymous):

3x - 5y + 2z = 7 4x + 9y - z = 2 x - 3y + 3z = 7 -------------- 12x+135y+1z= ranga it would be like this and what about the once of the = u multiple too or just +

OpenStudy (anonymous):

7+2+7=16 or 7*2*7=98

OpenStudy (anonymous):

ranga